# Phase diagram of the hard-sphere potential model in three and four dimensions using a pseudo-hard-sphere potential

### v1

@@ III.2 Integral equations theory
where $b(r)$ is the so-called bridge function [2]. In general, $b(r)$ is unknown and depends on the nature of the interaction potential. Although it is common to express $b(r)$ as an infinite series without a closed-form expression, several closed approximations for $b(r)$ have been developed and can be used to accurately study the thermodynamic properties of hard-sphere-like fluids. Several approximations for the bridge function can be found in the literature [39, 40], however in this work we focus on a modification to the semiphenomenological Verlet closure (MV) proposed by Kinoshita [41],
@@
which has been recently used to account for the structure, thermodynamics and depletion forces of binary mixtures of hard spheres even near thermodynamic instabilities, see, e.g., Ref. [42] and references therein.
The absolute value in Eq. (10) prevents divergence when the quantity $1+0.8\gamma(r)$ approaches zero, threby enhancing the numerical stability of the approximation [43]. A recent study [44] demonstrates that the MV closure more accurately reproduces the value of the radial distribution function at contact in a hard-sphere system compared to other bridge functions. However, to the best of our knowledge, Eq. (10) has not been used to solve the OZ equation (7) in four dimensions. Furthermore, the Percus–Yevick (PY) approximation, which provides an exact solution to the OZ equation in three dimensions [2], has historically been used to compute the equation of state for hard-sphere-like fluids [45]. Therefore, to ensure completeness and to showcase the accuracy of $b_{MV}(r)$ when applied to both hard and pseudo-hard systems, we also utilized the PY approximation [2], which is given by
@@
To numerically solve the OZ equation, we used a recently developed software written in the Julia programming language[46, 44]. This open-source code was designed to solve the OZ equation in arbitrary spatial dimensions and contains the necessary functions for calculating thermodynamic properties, such as isothermal compressibility and virial pressure. We obtain the radial distribution function for different packing fraction values in order to construct the equation of state. The OZ integral equation is numerically solved at the cutoff radius of $r_{c}=10\sigma$, we use $\mathrm{M}=2^{16}$ points within the interval of integration, and a grid spacing of $\mathrm{dr}=r_{c}/\mathrm{M}$. To reach the target densities, a density ramp is employed, and a sequence of initial values is used to create a good initial estimate, this is the Ng iteration scheme [38].
@@
Once we find the solution of the OZ equation, the corresponding $g(r)$ is used to construct the equation of state. The compressibility factor in $d$ dimensions can be written as follows [28],
[…]
where $d\mathbf{r}$ stands for the $d$-dimensional differential volume element. Although this relation involves an improper integral, the product of the $g(r)$ and the radial derivative of the $u_{WCA}(r)$ continuous potential is nonzero only over a very narrow interval. Since the potential decays rapidly at contact and quickly approaches zero at larger distances, and because $g(r)$ becomes nonzero only just before contact, the integration window in Eq. (13) is very narrow. Consequently, the grid used to solve the OZ equation must be sufficiently fine, containing enough points to accurately perform the numerical integration required to compute the compressibility factor via Eq. (13).
IV Fluid regime: Integral equations predictions and mean-field approximations
In this section we show the results of solving the OZ equation using the the MV closure, Eq. (10) and the PY closure, Eq. (11). The solutions are compared directly to the simulation results obtained for the BD simulations of the pHS model and the EDMD simulations of the hard-sphere model, in both $3D$ and $4D$. Additionally, we compare the results for the latter with a well-known mean-field approach.
@@ IV.1 Three dimensional equation of state
In Fig. 1, the equation of state found from solving the OZ is presented, along with the simulation results for both the hard-sphere and pseudo hard-sphere interaction potentials.
@@
From Fig. 1a), we observe that the PY closure is not precise enough to reproduce the compressibility factor at high filling fractions $\eta>0.4$, which has been a well-known fact from liquid state theory [2]. Furthermore, the difference between the HS and pHS models is not noticeable, alluding to the fact that both models are of comparable precision in the values obtained for the compressibility, something also observed in the simulations results.
@@
The deviation in the compressibility shows a clearer picture on the precision of the closures for the OZ equation by comparing it to the simulation results from BD simulations. In Fig. 1b), we observe that the deviations start to increase consistently starting for the PY closure from a small filling fraction of about $\eta\approx 0.2$, for both the HS and pHS models, indicating what we saw before that the PY is not a good closure for the HS model or even the pHS model. This is well-established, but it is good to see that we can reproduce here as well, even for the pHS model. On the other hand, the MV closure is highly accurate and reproduces the compressibility factor from simulations quite well, with the HS model having a smaller error in the measurement. This goes on to show that the closure remains a good approximation for solving the HS model, or similar interaction potentials, and it is a good idea to use this as a first attempt at solving the OZ equation for similar interaction potentials if no other knowledge of the closure relation is known beforehand.
@@ IV.2 Four dimensional equation of state
Now we turn our attention to the four dimensional case and the solution of the OZ equation, shown in Fig. 2. The case of four dimensions is interesting because there is a scarcity in the literature on this topic, although the topic of higher dimensional hard sphere models is of relevance, as has been pointed out previously. In particular, equations of state for mixtures of hard hyperspheres have been proposed and analyzed in four and five dimensions [50, 51, 52, 53, 54]. These mean-field approximations are written in terms of the equation of state of the one-component hard-hypersphere fluid. Therefore, an accurate expression for the compressibility factor of the one-component fluid is needed. A mean-field equation of state of hard $d$-dimensional hyperspheres is defined by Luban and Michels [55] as follows,
@@
which incorporates the exact expressions for the reduced virial coefficients $b_{2}$, $b_{3}$ and $b_{4}.$ The coefficients of the linear function $\zeta(\eta)=\zeta_{0}+\zeta_{1}(\eta/\eta_{cp})$, with $\eta_{cp}$ representing the crystalline close-packing value, are obtained using computer simulations and the known virial coefficients. All quantities in Eq. (14) can be explicitly found in Table 1.
@@
Table 1: Reduced virial coefficients and parameters of the Luban-Michels equation of state (14) for $d=4$ obtained from López de Haro, Santos, and B. Yuste [54].
@@
Panel (a) in Fig.2 shows that the Luban–Michels (LM) equation of state, defined in Eq.(14), exhibits excellent accuracy compared to simulation data. This remarkable agreement arises from the theoretical formulation of the equation as a ratio of polynomials that incorporates the virial coefficients, thereby yielding an accurate representation of the virial expansion across all densities up to the liquid–solid transition. Notably, it is surprising that the original work [29], which was based on a relatively small system of approximately $N=684$ particles, achieved sufficient precision to reproduce the compressibility factor observed in simulations of the much larger systems simulated in this work. This suggests that even with higher-accuracy data, further improvements in the predictive capability of the LM equation of state for the pressure of the four-dimensional hard-sphere fluid would likely be marginal.
@@
Figure 2a) shows that the PY closure performs significantly worse at higher filling fractions $\eta\geq 0.15$, where the compressibility factor begins to deviate from simulation results. In Figure 2b, deviations are observed starting at approximately $\eta\geq 0.15$, and they increase steadily with increasing filling fraction until a maximum deviation is reached, corresponding to the fluid-solid transition (data not shown). The fluid-solid transition is known to be challenging for the OZ equation due to the numerical instability of conventional algorithms. Thus, the PY closure is less reliable as an initial approach for obtaining thermodynamic properties of higher-dimensional hard-sphere-like fluids.
@@
On the other hand, the MV closure again exhibits good accuracy, as shown in Fig. 2a). However, at high densities, around $\eta\approx 0.2$, it fails to provide accurate estimates of the compressibility. This discrepancy is evident in Fig.2b), where the deviations increase with the filling fraction. Compared to the three-dimensional fluid, the deviations in the current case are larger and exhibit a steeper increase with density. This observation suggests that the original Verlet closure, defined only for the three-dimensional hard-sphere fluid [56], might require modifications in its functional form to properly account for the virial coefficients in higher dimensions. A procedure similar to that employed in Bedolla, Padierna, and Castañeda-Priego [43], which utilizes a machine learning framework to obtain the coefficients for a generalized functional MV closure, could be beneficial. Nevertheless, the MV closure demonstrates sufficient accuracy to reproduce the compressibility values observed in simulations. This promising result underscores the robustness of the MV closure in contexts beyond its original intended application. Future investigations should examine whether this closure remains effective in higher dimensions and determine its corresponding accuracy in those cases.
V Phase diagram of the pseudo hard-sphere potential
V.1 Three dimensional equation of state
@@
We present our results for the equation of state of the three dimensional pHS and the hard-sphere model. For the fluid branch, we simulate densities in the range of $\eta\in[0.01,0.49]$, and for the solid branch we simulate densities in the range $\eta\in[0.5,0.62]$, just below the value of the close packing of $\eta=0.65896$ [57, 58]. In Fig. 3a) we present the results obtained from simulations and we compare it with several mean-field equations of state, as well as those fitted with simulation data. We use the well-known Carnahan-Starling (CS) equation of state [2], and a re-fitted version of this equation of state using higher order virial coefficient due to Liu [59]. For the simulation parametrized equations of state, we use the virial coefficient fits from [12] for both the fluid and solid branches.
@@
Figure 3a) demonstrates that the pHS model accurately reproduces the qualitative behavior of the compressibility factor for the hard-sphere model, as well as the trends observed in mean-field equations of state. This outcome is consistent with prior findings, as the pHS model is well-known for its ability to replicate the equation of state [19] and the structural properties of the hard-sphere system [21, 22]. To quantitatively assess the accuracy of the pHS model, Figure 3b) presents the relative difference between the BD simulation results of the pHS model and various equations of state as a function of packing fraction. At higher packing fractions, the difference between the pHS and Carnahan-Starling (CS) equations of state is notably larger, which is expected since the CS equation of state is known to exhibit reduced precision in this regime. By contrast, the Liu equation of state aligns more closely with the simulation data, a result attributed to its inclusion of higher-order virial coefficients, which improve the accuracy of its estimation of the compressibility. The results obtained from EDMD simulations and the mKLM equation of state are nearly indistinguishable, as the mKLM equation is directly derived from EDMD simulation data. Consequently, they are compared equivalently, revealing that the pHS model performs well overall, except at high densities within the range $\eta\in[0.4,0.5]$. However, the relative difference in this region is smaller than previously reported [19], a result we attribute to the use of a larger number of particles, which improves the accuracy of pressure measurements.
@@
The coexistence pressure of the pHS model was measured to be $\beta P\sigma^{3}=11.66\pm 0.04$, a result that is in excellent agreement with the previously reported value for the pHS model by Jover et al. [18], $\beta P\sigma^{3}=11.65\pm 0.01$ [60]. In the study by Espinosa et al. [60], $NPT$ simulations were employed to determine the coexistence pressure of the pHS model at a reduced temperature of $k_{B}T/\epsilon=1.5$, following the methodology of Jover et al. [18]. These results confirm the suitability of the pHS model for determining coexistence properties and possibly investigating crystal nucleation. Notably, both measurements are consistent with the coexistence pressure of the hard-sphere model, reported as $\beta P\sigma^{3}=11.5645\pm 0.0005$ [34]. However, given the relatively higher uncertainty in the pHS measurement, it remains an open question whether the free energy difference between the face-centered cubic (FCC) lattice and the hexagonal close-packed (HCP) structure [5] can be distinguished using the pHS model.
V.2 Four dimensional equation of state

### v2

@@ III.2 Integral equation theory
where $b(r)$ is the so-called bridge function [2]. In general, $b(r)$ is unknown and depends on the nature of the interaction potential. Although it is common to express $b(r)$ as an infinite series without a closed-form expression, several closed approximations for $b(r)$ have been developed and can be used to accurately study the thermodynamic properties of HS-like fluids. Several approximations for the bridge function can be found in the literature [40, 41], however, in this work we focus on a modification to the semiphenomenological Verlet closure (MV) proposed by Kinoshita [42],
@@
which has been recently used to account for the structure, thermodynamics, and depletion forces of binary mixtures of HS’s even near thermodynamic instabilities; see, e.g., [43] and references therein.
The absolute value in Eq. (11) prevents divergence when the quantity $1+0.8\gamma(r)$ approaches zero, thus improving the numerical stability of the approximation [44]. A recent study [45] demonstrates that the MV closure more accurately reproduces the value of the radial distribution function at contact in a HS system compared to other bridge functions. However, to our knowledge, Eq. (11) has not been used to solve the OZ equation (8) in four dimensions. Furthermore, the Percus–Yevick (PY) approximation, which provides an exact solution to the OZ equation in three dimensions [2], has historically been used to compute the equation of state for HS-like fluids [46]. Therefore, to ensure completeness and to show the precision of $b_{MV}(r)$ when applied to both hard and pseudo-hard systems, we also utilized the PY approximation [2], which is given by
@@
To numerically solve the OZ equation, we used a recently developed software written in the Julia programming language[47, 45]. This open-source code was designed to solve the OZ equation in arbitrary spatial dimensions and contains the necessary functions for calculating thermodynamic properties, such as the isothermal compressibility and the virial pressure. We obtain the radial distribution function for different packing fraction values in order to construct the equation of state. The OZ integral equation is numerically solved at the cutoff radius of $r_{c}=10\sigma$, we use $\mathrm{M}=2^{16}$ points within the interval of integration, and a grid spacing of $\mathrm{dr}=r_{c}/\mathrm{M}$. To reach the target densities, a density ramp is employed, and a sequence of initial values is used to create a good initial estimate, this is the Ng iteration scheme [39].
@@
Once we find the solution of the OZ equation, the corresponding $g(r)$ is used to construct the equation of state using Eq. (5). The latter involves an improper integral of the product of the $g(r)$ and the radial derivative of the continuous potential $u_{pHS}(r)$. Since the potential decays rapidly at contact and quickly approaches zero at larger distances, and because $g(r)$ becomes nonzero only just before contact, the integration window in Eq. (5) is very narrow. Consequently, the grid used to solve the OZ equation must be sufficiently fine, containing enough points to accurately perform the numerical integration required to compute the compressibility factor via Eq. (5).
IV Fluid regime: Integral equation predictions and mean-field approximations
In this section we show the results of solving the OZ equation using the the MV closure, Eq. (11) and the PY closure, Eq. (12). The solutions are compared directly to the simulation results obtained for the BD simulations of the pHS model and the EDMD simulations of the HS model, in both $3D$ and $4D$. Additionally, we compare the results for the latter with a well-known mean-field approach.
@@ IV.1 Three dimensional equation of state
In Fig. 1, the equation of state found from solving the OZ is presented, along with the simulation results for both the HS and pHS interaction potentials.
@@
From Fig. 1a), we observe that the PY closure is not precise enough to reproduce the compressibility factor at high filling fractions $\eta>0.4$, which has been a well-known fact from liquid state theory [2]. Furthermore, the difference between the HS and pHS models is not noticeable, alluding to the fact that both models are of comparable precision in the values obtained for the compressibility factor, something also observed in the simulations results.
@@
The deviation in the compressibility factor shows a clearer picture on the precision of the closures for the OZ equation by comparing it to the simulation results from BD simulations. In Fig. 1b), we observe that the deviations start to increase consistently starting for the PY closure from a small filling fraction of about $\eta\approx 0.2$, for both the HS and pHS models, indicating what we saw before that the PY is not a good closure for the HS model or even the pHS model. This is well-established, but it is good to see that we can reproduce here as well, even for the pHS model. On the other hand, the MV closure is highly accurate and reproduces the compressibility factor from simulations quite well, with the HS model having a smaller error in the measurement. This goes on to show that the closure remains a good approximation for solving the HS model, or similar interaction potentials, and it is a good idea to use this as a first attempt at solving the OZ equation for similar interaction potentials if no other knowledge of the closure relation is known beforehand.
@@ IV.2 Four dimensional equation of state
Now we turn our attention to the four dimensional case and the solution of the OZ equation, shown in Fig. 2. The case of four dimensions is interesting because there is a scarcity in the literature on this topic, although the topic of higher dimensional HS models is of relevance, as has been pointed out previously. In particular, equations of state for mixtures of hard hyperspheres have been proposed and analyzed in four and five dimensions [51, 52, 53, 54, 55]. These mean-field approximations are written in terms of the equation of state of the one-component hard-hypersphere fluid. Therefore, an accurate expression for the compressibility factor of the one-component fluid is needed. A mean-field equation of state of hard $d$-dimensional hyperspheres is defined by Luban and Michels [56] as follows,
@@
which incorporates the exact expressions for the reduced virial coefficients $b_{2}$, $b_{3}$ and $b_{4}.$ The $k$-th reduced virial coefficient can be written in terms of $V_{d}$[57]
$b_{k}=\left(\frac{V_{d}}{2^{d}}\sigma^{d}\right)^{-(k-1)}B_{k}\,,$ (15)
with $B_{k}$ being the $k$-th virial coefficient. The coefficients of the linear function $\zeta(\eta)=\zeta_{0}+\zeta_{1}(\eta/\eta_{cp})$, with $\eta_{cp}$ representing the crystalline close-packing value, are obtained using computer simulations and the known virial coefficients. All quantities in Eq. (14) can be found explicitly in Table 1.
@@
Table 1: Reduced virial coefficients and parameters of the Luban-Michels equation of state (14) for $d=4$ obtained from López de Haro et al. [55]
@@
Panel (a) in Fig. 2 shows that the Luban–Michels (LM) equation of state, defined in Eq. (14), exhibits excellent accuracy compared to simulation data. This remarkable agreement arises from the theoretical formulation of the equation as a ratio of polynomials that incorporates the virial coefficients, thereby yielding an accurate representation of the virial expansion across all densities up to the liquid–solid transition. Notably, it is surprising that the original work [30], which was based on a relatively small system of approximately $N=684$ particles, achieved sufficient precision to reproduce the compressibility factor observed in simulations of the much larger systems simulated in this work. This suggests that even with higher-accuracy data, further improvements in the predictive capability of the LM equation of state for the pressure of the four-dimensional HS fluid would likely be marginal.
@@
Figure 2a) shows that the PY closure performs significantly worse at higher filling fractions $\eta\geq 0.15$, where the compressibility factor begins to deviate from simulation results. In Figure 2b, deviations are observed starting at approximately $\eta\geq 0.15$, and they increase steadily with increasing filling fraction until a maximum deviation is reached, corresponding to the fluid-solid transition (data not shown). The fluid-solid transition is known to be challenging for the OZ equation due to the numerical instability of conventional algorithms. Thus, the PY closure is less reliable as an initial approach for obtaining thermodynamic properties of higher-dimensional HS-like fluids.
@@
On the other hand, the MV closure again exhibits good accuracy, as shown in Fig. 2a). However, at high densities, around $\eta\approx 0.2$, it fails to provide accurate estimates of the compressibility factor. This discrepancy is evident in Fig. 2b), where the deviations increase with the filling fraction. Compared to the three-dimensional fluid, the deviations in the current case are larger and exhibit a steeper increase with density. This observation suggests that the original Verlet closure, defined only for the three-dimensional HS fluid [58], might require modifications in its functional form to properly account for the virial coefficients in higher dimensions. A procedure similar to that employed in Bedolla, Padierna, and Castañeda-Priego [44], which utilizes a machine learning framework to obtain the coefficients for a generalized functional MV closure, could be beneficial. Nevertheless, the MV closure demonstrates sufficient accuracy to reproduce the compressibility factor values observed in simulations. This promising result underscores the robustness of the MV closure in contexts beyond its original intended application. Future investigations should examine whether this closure remains effective in higher dimensions and determine its corresponding accuracy in those cases.
V Long-time self-diffusion coefficient of the pseudo hard-sphere potential
Although this contribution mainly deals with the equilibrium phase diagram and, therefore, thermodynamic properties are determined, we consider it relevant to analyze and discuss whether the particle trajectories obtained from the Brownian dynamics (BD) simulations of the pHS potential correctly reproduce the transport properties of hard spheres in the so-called diffusive regime. This case corresponds to the dynamics of a colloidal dispersion made up of hard spheres [59]. However, as we discuss below, there exists a heuristic criterion based on the long-time self-diffusion coefficient that also provides an estimation of the liquid-solid transition [60]. This route will be an independent confirmation that the pHS model reproduces not only the thermodynamic properties of hard spheres but also the transport phenomena.
As explained above, we have extracted the ratio $D_{L}/D_{0}$ from the mean-square displacement. The latter is computed using the pHS potential for every packing fraction. Then, in Fig. 3, we present the BD simulation results without hydrodynamic interactions (HI), indicated by square symbols. These results are compared with experimental measurements of colloidal suspensions [61] nearly interacting as hard spheres, labeled as EXP. We also include results from dynamic Monte Carlo simulations of hard spheres without HI [62], as well as the theoretical predictions of Medina- Noyola with and without HI [59]; this dynamical approach made use of the Percus-Yevick approximation as the static input. We first compare the predictions of the pHS potential with the experiments. As seen, the agreement is good at low volume fractions ($\eta<0.2$); however, the discrepancies become more pronounced at higher densities. We attribute this deviation to the presence of HI in the experimental systems. In fact, the original experimental work includes a comparison with the simulation data that accounts for HI [62], highlighting significant differences at higher densities. Interestingly, the theoretical framework including HI seems to perform well at high concentrations, but the case without HI nicely follows the predictions from the pHS, which basically reproduce the dynamic Monte Carlo data for hard spheres. This level of agreement allows us to conclude that the Brownian dynamics of the pHS potential correctly reproduces the diffusive behavior of hard spheres without HI. Nonetheless, results from quasi-two-dimensional hard-disk experiments [63] suggest that HI have a negligible impact on the long-time self-diffusion coefficient, with good agreement observed between experimental and simulation data. This raises the question of whether a similar dynamical behavior might hold in three-dimensional systems. We propose that a renewed experimental effort to measure the long-time self-diffusion coefficient together with the use of the pHS potential in Brownian dynamics simulations including explicitly HI could help resolve this discrepancy and clarify the role of HI in dense colloidal suspensions; work along this line is in progress.
Figure 3: The reduced long-time self-diffusion coefficient, $D_{L}/D_{0}$, as a function of the packing fraction in three dimensions, $\eta=\pi\rho\sigma^{3}/6$, is shown. The Brownian dynamics simulations of the pHS potential are shown (close square) with error bars smaller than than the symbol size. The results are compared against experiments (close diamonds) of colloidal suspensions (EXP) [61] and dynamic Monte Carlo simulations (closed stars) of hard spheres without hydrodynamic interactions [62] (stars; MC-BD). Also, theoretical predictions from Medina-Noyola [59] (lines; Theory) with and without HI are also displayed.
Furthermore, we note that the BD results with the pHS potential also show excellent agreement with the Lowen’s freezing criterion[60], since the value of the long-time self-diffusion coefficient at a packing fraction of $\eta=0.48$ is $D_{L}/D_{0}=0.143\pm 0.001$, which is close to the predicted value of $D_{L}/D_{0}=0.1$ and this case corresponds to the liquid-solid transition point for hard spheres, $\eta=0.49$, which we will study in detail in the following section.
VI Phase diagram of the pseudo hard-sphere potential
VI.1 Three dimensional equation of state
@@
We present our results for the equation of state of the three dimensional pHS and the HS model. For the fluid branch, we simulate densities in the range of $\eta\in[0.01,0.49]$, and for the solid branch, we simulate densities in the range $\eta\in[0.5,0.62]$, just below the value of the densest packing $\eta_{CP}=\pi/\sqrt{18}$. In Fig. 4a) we present the results obtained from simulations and we compare it with several mean-field equations of state, as well as those fitted with simulation data. We use the well-known Carnahan-Starling (CS) equation of state [2], and a re-fitted version of this equation of state using higher order virial coefficient due to Liu [64]. For the simulation parametrized equations of state, we use the virial coefficient fits from [12] for both the fluid and solid branches.
@@
Figure 4a) demonstrates that the pHS model accurately reproduces the qualitative behavior of the compressibility factor for the HS model, as well as the trends observed in mean-field equations of state. This outcome is consistent with prior findings, as the pHS model is well-known for its ability to replicate the equation of state [20] and the structural properties of the HS system [22, 23]. To quantitatively assess the accuracy of the pHS model, Figure 4b) presents the relative difference between the BD simulation results of the pHS model and various equations of state as a function of packing fraction. At higher packing fractions, the difference between the pHS and Carnahan-Starling (CS) equations of state is notably larger, which is expected since the CS equation of state is known to exhibit reduced precision in this regime. By contrast, the Liu equation of state aligns more closely with the simulation data, a result attributed to its inclusion of higher-order virial coefficients, which improve the accuracy of its estimation of the compressibility factor. The results obtained from EDMD simulations and the mKLM equation of state are nearly indistinguishable, as the mKLM equation is directly derived from EDMD simulation data. Consequently, they are compared equivalently, revealing that the pHS model performs well overall, except at high densities within the range $\eta\in[0.4,0.5]$. However, the relative difference in this region is smaller than previously reported [20], a result we attribute to the use of a larger number of particles, which improves the accuracy of pressure measurements.
@@
The coexistence pressure of the pHS model was measured to be $\beta P\sigma^{3}=11.66\pm 0.04$, a result that is in excellent agreement with the previously reported value for the pHS model by Jover et al. [19], $\beta P\sigma^{3}=11.65\pm 0.01$ [65]. In the study by Espinosa et al. [65], $NPT$ simulations were employed to determine the coexistence pressure of the pHS model at a reduced temperature of $k_{B}T/\epsilon=1.5$, following the methodology of Jover et al. [19]. These results confirm the suitability of the pHS model for determining coexistence properties and possibly investigating crystal nucleation. Notably, both measurements are consistent with the coexistence pressure of the HS model, reported as $\beta P\sigma^{3}=11.5645\pm 0.0005$ [35]. However, given the relatively higher uncertainty in the pHS measurement, it remains an open question whether the free energy difference between the face-centered cubic (FCC) lattice and the hexagonal close-packed (HCP) structure [5] can be distinguished using the pHS model.
VI.2 Four dimensional equation of state